# Continuous-variable quantum key distribution with noisy squeezed states

### v1

@@ Abstract
We address the role of noisy squeezing in security and performance of continuous-variable (CV) quantum key distribution (QKD) protocols. Squeezing has long been recognized for its numerous advantages in CV QKD, such as enhanced robustness against channel noise and loss, and improved secret key rates. However, the noise of the squeezed states, that unavoidably originates already from optical loss in the source, raises concerns about its potential exploitation by an eavesdropper. This is particularly relevant if this noise is pessimistically assumed untrusted. We address the allocation of untrusted noise within a squeezed state and show that anti-squeezing noise is typically more harmful for security of the protocols, as it potentially provides more information to an eavesdropper. Although the anti-squeezing noise may not directly contribute to the generated key data, it is involved in parameter estimation and can in fact be harmful even if considered trusted. Our study covers the effects of anti-squeezing noise in both the asymptotic and finite-size regimes. We highlight the positive effects and limitations of imposing trust assumption on anti-squeezing noise. Additionally, we emphasize the detrimental impact of untrusted noise in both fiber and free-space fading links. Our findings offer essential insights for practical implementations and optimization of squeezed-state CV QKD protocols in realistic scenarios.
@@ 1 Introduction
Quantum key distribution (QKD) [1, 2, 3, 4, 5] is arguably the most mature of the quantum technologies, which has its goal in the development of protocols for provably secure distribution of symmetrical cryptographic keys. Recent advances in the field are particularly concerned with the continuous-variable (CV) [6] approach to QKD, focused on realizations using off-the-shelf high-speed and efficient coherent homodyne detection. CV QKD was intensively developed, studied, and tested using coherent signal states [7, 8, 9, 10, 11], typically using Gaussian quadrature modulation, compatible with the most advanced security proofs [12, 13]. Alternatively, CV QKD can be realized using quadrature modulation of squeezed states, which was the initial proposal for Gaussian CV QKD [14]. Such states are known to have quantum fluctuations of a quadrature observable of light field (referred to as a squeezed quadrature) to be suppressed under the level of vacuum fluctuations (being also known as the shot-noise level) [15]. To comply with the Heisenberg uncertainty principle, a complementary quadrature observable then has the fluctuations above the shot-noise level, and is referred to as the anti-squeezed (AS) quadrature. In the ideal case of a pure squeezed state, the AS quadrature variance is inverse to the squeezed one.
@@
Importantly, squeezed states are never perfectly pure in practice and contain some excess noise in the AS quadrature above the level of inverse of the squeezing variance (currently the highest achieved purity of strongly squeezed state is 89% [23]). This is concerned with the fact that optical loss is unavoidably present in a device that produces squeezed states. Indeed, in the quantum description, losses linearly couple a squeezed state to the vacuum one. The squeezed quadrature variance, being below the shot noise, then becomes larger (hence corresponding to the lower level of squeezing). On the other hand, the AS quadrature variance, being much above the shot noise, becomes reduced, but is still larger than the inverse of the squeezed variance after the losses. This leads to effective emergence of the AS excess noise, which is larger, the stronger are the initial squeezing and the optical losses in the source. The typical values of AS noise observed in the experiment can range from 0.8dB [24], to 4.3dB and 8.4dB [25, 26]. Such noise was recently studied in the context of phase sensitive amplification in optical communication in Ref.[27].
@@
The AS noise is typically assumed trusted in CV QKD security analysis, being a part of a trusted sender station [17, 28, 19] and hence does not directly affect the security of CV QKD, at least in the idealized case of infinitely large data ensemble (also referred to as asymptotic case), which implies perfect parameter estimation in CV QKD. Furthermore, the outcomes of the measurement of the AS quadrature do not directly contribute to the key data. This is the main reason why the role of AS noise was not previously studied in the context of CV QKD. However, an eavesdropper may at least partially control the preparation device and hold the purification of the AS noise, which undermines the trust assumption on this type of noise. Also, the trusted parties may not be able to distinguish the AS noise from the noise introduced in the channel, and hence assume the noise to be untrusted. Furthermore, in the finite-size setting of a limited data ensemble, the parameter estimation, which is an essential part of a CV QKD protocol, may probabilistically rely on the measurements of the AS quadrature and be corrupt even by the trusted AS noise. Finally, the AS noise, even if assumed trusted, leads to untrusted excess noise in the fluctuating free-space channels and imposes limitations on the levels of squeezing already for the pure states [21]. Hence it is necessary to analyse the role of AS noise in CV QKD, which we do in the current paper. We first show how the noise in the squeezed states appears already due to optical attenuation and identify the most typically pessimistic case for CV QKD to assume that all the untrusted noise is allocated in the AS quadrature. We develop a model for incorporating the AS noise in security analysis of CV QKD. Using the model, we show certain positive effects of trusted AS noise in CV QKD in the asymptotic regime of infinitely large data ensembles. On the other hand, we show the limitations imposed on the secure range of the protocols if the assumption on the trusted AS noise is waived. We also present the negative effects on the parameter estimation and subsequent reduction of security range if the AS quadrature measurements are used for parameter estimation, which holds true already for the trusted AS and is worsen if the AS excess noise is untrusted. Finally, we analyse the negative role of AS noise in fluctuating free-space channels in both trusted and untrusted scenarios. The obtained results are essential for realization of the squeezed-state CV QKD protocols in practical scenarios.
@@
Our paper is structured as follows: we start with the simplest description of noisy squeezed states as the result of optical losses in Sec. 2, and adopt the most conservative and general approach in Sec. 3 where we describe the protocol, build the model for purification of trusted AS noise, and derive the relevant entropic quantities; in Sec. 4.1 we discuss the effects of trusted and untrusted AS in the asymptotic regime; in Sec. 4.2 we address the role of AS in the parameter estimation and obtain the implications on the key rate and security range of the protocol as well as suggest the optimal protocol implementations in the presence of AS noise; in Sec 5 we study the role of AS in fluctuating free space channels, and then finish the paper with concluding remarks.
2 Noisy squeezing due to optical losses
Squeezed states, typically generated using optical parametric oscillators [15], are unavoidably noisy already due to the optical losses in the cavity and inefficient collection of light at the cavity output. A single-mode pure squeezed state, described by the squeezed and AS variances $V_{S},1/V_{S}$, undergoing losses can be modeled as coupling to a vacuum mode on a beamsplitter with transmittance $\eta$. The resulting state after the beamsplitter interaction [29] is described by the quadrature variances $\eta V_{S}+1-\eta,\eta/V_{S}+1-\eta$ and contains the fraction of vacuum noise (in shot-noise units, SNU), scaled by the beamsplitter reflectance $1-\eta$. The Gaussian purity of the state then becomes $\mu=\left[1+(\eta-\eta^{2})(1/V+V-2)\right]^{-1/2}$ and is minimized at $\eta=1/2$. When detected, without precise knowledge of coupling loss $\eta$ and initial squeezing $V_{S}$, one cannot pinpoint the exact distribution of added noise ($\epsilon_{x}$ and $\epsilon_{p}$) in each quadrature. In the context of CV QKD, when the actual $V_{S}$ and $\eta$ are not known, it can be assumed that the resulting state contains a pure squeezed state with variances $V,1/V$ and some excess noise $\epsilon_{x},\epsilon_{p}$ in either of the quadratures. The allocation of the noise once it is trusted (not controlled by Eve) does not influence the security of CV QKD. However, when the noise in the squeezed signal states is not trusted (controlled by Eve), its allocation to particular quadratures has direct security ramifications and we must adopt the most conservative stance. The latter typically implies admitting the largest quantitative amount of noise, i.e. treating all the noise as contained in the AS quadrature, see Appendix A for further details.
@@ 3 CV QKD protocol with AS excess noise
To investigate the role of AS noise in the squeezed-state CV-QKD, we consider the prepare-and-measure (P&M) squeezed-state protocol, as shown in Fig. 2, in which a trusted party (hereafter referred to as Alice) generates a squeezed signal state with squeezed quadrature variance V and AS quadrature variance $1/V+\Delta V$, $\Delta V$ being the excess AS noise. Alice then modulates the squeezed quadrature and AS quadrature (here and further with no loss of generality we assume the $x$-quadrature to be squeezed) according to Gaussian distributed random variables $x_{m}$ and $p_{m}$ with variances $V_{x}$ and $V_{p}$, respectively. Note that while the secret key is typically obtained from the measurements of the squeezed quadrature, we assume that both quadratures are modulated in the general case, particularly for estimating the channel parameters. The modulated squeezed state is sent to a remote trusted party (hereafter referred to as Bob) via a quantum channel characterized by transmittance $T$, and quadrature excess noise $\epsilon$ (which we, without loss of generality, quantify with respect to the channel input). Bob then performs a homodyne detection to measure either x- or p-quadrature of the quantum state he receives. Alternatively, Bob can measure both quadratures simultaneously with a generally unbalanced heterodyne detector (double homodyne) with the balancing $t_{het}$ as shown in Fig. 2b. After a sufficient number of transmitted pulses are measured, trusted parties perform error correction and privacy amplification on their data to obtain the secret key. The lower bound on the secret key, secure against collective attacks (directly extendable to security against general attacks in the asymptotic regime [13]) is given by:
@@
Figure 2: a) Prepare and measure scheme of a CV QKD protocol with noisy squeezed states. Alice modulates the squeezed state, characterized by the squeezed variance $V$ and AS excess noise $\Delta V$, according to normally distributed random variables and sends the Gaussian modulated squeezed state through a quantum channel characterized by transmittance $T$ and excess noise $\epsilon$. Bob at his end performs homodyne detection or b) unbalanced heterodyne detection with $t_{het}$. c) Equivalent general source model based on the balanced coupling of two oppositely squeezed states in modes $A_{1}$ and $B$ with different squeezed variances $V_{1}$, $V_{2}$ and a balanced heterodyne measurement at the Alice’s side, fed by a squeezed state with variance $V_{m}$ in mode $A_{2}$.
@@
The trusted parties may assume the AS noise $\Delta V$ to be either trusted (hence, being out of control by an eavesdropper) or untrusted (fully controlled by Eve). We further denote the trusted and untrusted AS noise as $\Delta V_{t}$ and $\Delta V_{u}$, respectively. We model the trusted AS noise by redefining the modulation variance of the AS quadrature $V_{p}\rightarrow V_{p}+\Delta V_{t}$. On the other hand, the untrusted AS noise $\Delta V_{u}$ is added to the AS quadrature similarly to the channel noise $\epsilon$.
@@
To study the optimal performance of CV QKD with untrusted AS noise (on top of symmetric noise $\epsilon$) in the asymptotic regime, we only consider the measurement of squeezed quadrature by Bob.
@@ 4.1 Asymptotic key rate
In the presence of channel noise, no simple analytical expressions can be obtained for the Holevo bound, as we rely on the 3-mode purification described in Sec.3. The presence of trusted noise can improve tolerance to untrusted channel noise, as can be seen from the plots in Fig. 3, where we model the performance of the protocol assuming realistic error correction efficiency of $\beta=0.95$ [37] and optimizing the modulation variance. The improvement by the trusted AS noise is more evident in DR than in RR, being a case of the ”fighting noise with noise” method [16, 36]. Interestingly, a moderately squeezed source ($V=0.5$ SNU) in the presence of optimal trusted AS noise has a higher tolerance to channel noise than the protocol with a higher degree of pure signal squeezing of the source state. On the other hand, when the AS noise cannot be trusted, the tolerance of the protocol to channel noise is significantly reduced, as can be seen from the plots in Fig. 4, as the presence of untrusted AS noise increases the Holevo bound. This emphasises the need to control and to estimate noise in the AS quadrature, which would otherwise lead to underestimation of the Holevo bound.
@@ 4.2 Finite-size key rate
A necessary requirement to verify the security of QKD is to estimate the channel parameters from which the Holevo bound can be found. Channel parameters can only be estimated within a certain confidence interval, which is inversely propositional to the number of states sent to Bob. Typically part of the received data is used for parameter estimation, and the rest is used for the key generation [38]. It has been also shown that all the measurements can be used for both parameter estimation and key generation by performing error correction before parameter estimation [39]. For completeness, we consider both cases, when part of the measurements are disclosed for parameter estimation (see B), and when all the measurements are used for parameter estimation and key generation.
@@
In order not to underestimate the information available to the eavesdropper, we use the lower bound for the confidence interval of channel transmittance $T^{low}$ and the upper bound for channel noise $V_{\epsilon}^{up}$. We consider $\mathbf{E}(T^{low})=T-6.5\Upsilon$ and $\mathbf{E}(V_{\epsilon}^{up})=V_{\epsilon}+6.5s$ corresponding to the estimation error probability of $10^{-10}$, $\sigma$ and $s$ are the standard deviations of the channel transmittance and noise, respectively. In order to effectively use the available resources, we extend the parameter estimation from [12, 38] to use the measurements of both quadratures to potentially make a better estimation of the channel parameters (refer to B for details). We obtain the following finite-size key rate:
@@
In Fig. 5, our analysis reveals distinct trends based on the degree of squeezing in the source state. Imbalanced heterodyne measurements demonstrate superior performance with small block sizes ($N=10^{6}$) for highly squeezed states (dashed plots), while biased homodyne detection becomes more favorable with larger block sizes ($N=10^{7}$), see Fig. 54(b), 4(d) and 4(f). Notably, biased homodyne measurements consistently outperform other strategies for moderately squeezed states. In scenarios marked by elevated trusted noise and reduced block lengths (Fig. 54(e)), imbalanced heterodyne measurement proves significantly more efficacious than biased homodyne measurement. Furthermore, we present key rates when the AS quadrature measurement in imbalanced heterodyne measurements is disclosed publicly (red plots in Fig. 5 and 6), showcasing its potential contributions to parameter estimation and the selection of optimal error correction codes.
Fig. 6 adds complexity to the narrative, revealing that the efficacy of imbalanced heterodyne measurements is contingent on the presence of channel noise. Specifically, imbalanced heterodyne measurement emerges as consistently superior in scenarios characterized by a substantial amount of channel noise ($\epsilon>0.1$ SNU), while biased homodyne measurement excels in low-channel-noise conditions ($\epsilon<0.1$ SNU). The interplay of measurement strategies with channel noise highlights the nuanced considerations necessary for optimizing squeezed-state CV quantum communication systems.

### v2

@@ Abstract
We address the crucial role of noisy squeezing in security and performance of continuous-variable (CV) quantum key distribution (QKD) protocols. Squeezing has long been recognized for its numerous advantages in CV QKD, such as enhanced robustness against channel noise and loss, and improved secret key rates. However, the excess noise of the squeezed states, that unavoidably originates already from optical loss as well as other imperfections in the source, raises concerns about its potential exploitation by an eavesdropper. For the widely adopted trust assumption on the excess noise in the signal states, we confirm the stability of the protocol against the noisy squeezing in both purely attenuating as well as noisy channels in the asymptotic limit, which implies perfect parameter estimation. In the finite-size regime we show that this stability largely holds at up to $10^{7}$ data points using optimal biased homodyne detection for key distribution and parameter estimation. Untrusted assumption on the noisy squeezing, on the other hand, introduces additional security bounds on the squeezing excess noise already in the asymptotic regime, which is further enforced by the finite-size effects. Additionally, we show the critical negative role of noisy squeezing in the case of atmospheric free-space channels, already in the trusted-noise and asymptotic assumptions, which emphasizes the importance of squeezing purity in the free-space quantum channels. Our results pave the way towards practical realization of squeezed-state CV QKD protocols in both fibre and free-space channels.
@@ 1 Introduction
Quantum key distribution (QKD) [1, 2, 3, 4, 5] is arguably the most mature of the quantum technologies, which has its goal in the development of protocols for provably secure distribution of symmetrical cryptographic keys. Recent advances in the field are particularly concerned with the continuous-variable (CV) [6] approach to QKD, focused on realizations using off-the-shelf high-speed and efficient coherent homodyne detection. CV QKD was intensively developed, studied, and tested using coherent signal states [7, 8, 9], typically using Gaussian quadrature modulation, compatible with the most advanced security proofs [10, 11]. In recent years, significant progress has been made in enhancing the performance and the security of CV-QKD protocols under realistic conditions, including improved noise tolerance [12, 13], higher key generation speed [14, 15, 16], and longer secure distance [17, 18]. All of which can be further refined by integrating automated machine learning algorithms [19, 20].
Alternatively, CV QKD can be realized using quadrature modulation of squeezed states, which was the initial proposal for Gaussian CV QKD [21]. Such states are known to have quantum fluctuations of a quadrature observable of light field (referred to as a squeezed quadrature) to be suppressed under the level of vacuum fluctuations (being also known as the shot-noise level) [22]. To comply with the Heisenberg uncertainty principle, a complementary quadrature observable then has the fluctuations above the shot-noise level, and is referred to as the anti-squeezed (AS) quadrature. In the ideal case of a pure squeezed state, the AS quadrature variance is inverse to the squeezed one (putting =2Planck-constant-over-$\hbar=2$).
@@
Moreover, squeezed-state CV QKD provides certain advantages over alternative QKD protocols. Studies have demonstrated that in low phase noise but high thermal noise environments, squeezed-state CV QKD can tolerate more loss compared to DV QKD protocols like the six-state protocol [30]. Additionally, non-Gaussian CV QKD protocols, which integrate operations like photon subtraction, can extend the communication range, but their implementation complexity makes squeezed-state CV QKD a favorable choice for practical applications [31]. These distinctions underline the importance of assessing the security and performance of squeezed states under realistic conditions.
@@
Importantly, squeezed states are never perfectly pure in practice and contain some excess noise in the AS quadrature above the level of inverse of the squeezing variance (currently the highest achieved purity of strongly squeezed state is 89% [32]). This is concerned with the fact that optical loss, phase noise, parasitic nonlinearities etc. are unavoidably present in a device that produces squeezed states. This leads to effective emergence of the AS excess noise, which is larger, the stronger are the initial squeezing and the imperfections in the source. The typical values of AS noise observed in the experiment can range from 0.8dB [33], to 4.3dB and 8.4dB [34, 35]. Such noise was recently studied in the context of phase sensitive amplification in optical communication in Ref.[36].
The AS noise is typically assumed trusted in CV QKD security analysis, being a part of a trusted sender station [24, 37, 26, 38] and hence does not directly affect the security of CV QKD, at least in the idealized case of infinitely large data ensemble (also referred to as asymptotic case), which implies perfect parameter estimation in CV QKD. Furthermore, the outcomes of the measurement of the AS quadrature do not directly contribute to the key data. This is the main reason why the role of AS noise was not previously studied in the context of CV QKD.
Trusted noise refers to noise that arises from known, well-characterized imperfections within the sender’s and receiver’s own devices, located in their isolated stations. Because these imperfections are closely monitored and their effects are understood, this noise is assumed not to assist an eavesdropper in gaining information about the key. In contrast, untrusted noise is any noise that cannot be confidently attributed to such internal, controlled sources. It includes noise originating from external or poorly characterized components, and thus cannot be ruled out as a potential avenue for an eavesdropper’s attack.
@@
However, an eavesdropper may at least partially control the preparation device and hold the purification of the AS noise, which undermines the trust assumption on this type of noise. Also, the trusted parties may not be able to distinguish the AS noise from the noise introduced in the channel, and hence assume the noise to be untrusted. Furthermore, in the finite-size setting of a limited data ensemble, the parameter estimation, which is an essential part of a CV QKD protocol, may probabilistically rely on the measurements of the AS quadrature and be corrupt even by the trusted AS noise. Finally, the AS noise, even if assumed trusted, leads to untrusted excess noise in the fluctuating free-space channels and imposes limitations on the levels of squeezing already for the pure states [28].
Hence it is necessary to analyse the role of AS noise in CV QKD, which we do in the current paper. We develop a purification model for incorporating the AS noise in security analysis of CV QKD. Using the model, we demonstrate that trusted AS noise in CV-QKD has no negative impact and can even have a slight positive effect in the asymptotic regime of infinitely large data ensembles. However, when the assumption of trusted AS noise is waived, the protocol’s security is substantially and adversely impacted. We also present the negative effects on the parameter estimation and subsequent reduction of security range if the AS quadrature measurements are used for parameter estimation, which holds true already for the trusted AS and is worsen if the AS excess noise is untrusted. Finally, we analyse the negative role of AS noise in fluctuating free-space channels in both trusted and untrusted scenarios. The obtained results are essential for realization of the squeezed-state CV QKD protocols in practical scenarios and further development of the protocols with realistic squeezed states.
@@
Our paper is structured as follows: we describe the protocol and adopt the most conservative and general approach in Sec. 2 where we describe the protocol, build the model for purification of trusted AS noise, and derive the relevant entropic quantities; in Sec. 3.1 we discuss the effects of trusted and untrusted AS in the asymptotic regime; in Sec. 3.2 we address the role of AS in the parameter estimation and obtain the implications on the key rate and security range of the protocol as well as suggest the optimal protocol implementations in the presence of AS noise; in Sec 4 we study the role of AS in fluctuating free space channels, and then finish the paper with concluding remarks.
@@ 2 CV QKD protocol with AS excess noise
To investigate the role of AS noise in the squeezed-state CV-QKD, we consider the prepare-and-measure (P&M) squeezed-state protocol, as shown in Fig. 1, in which a trusted party (hereafter referred to as Alice) generates a squeezed signal state with squeezed quadrature variance V and AS quadrature variance $1/V+\Delta V$, $\Delta V$ being the excess AS noise. Alice then modulates the squeezed quadrature and AS quadrature (here and further with no loss of generality we assume the $x$-quadrature to be squeezed) according to Gaussian distributed random variables $x_{m}$ and $p_{m}$ with variances $V_{x}$ and $V_{p}$, respectively. Note that while the secret key is typically obtained from the measurements of the squeezed quadrature, we assume that both quadratures are modulated in the general case, particularly for estimating the channel parameters. State preparation imperfections (including modulation) are attributed to the squeezed light source. The modulated squeezed state is sent to a remote trusted party (hereafter referred to as Bob) via a quantum channel characterized by transmittance $T$, and quadrature excess noise $\epsilon$ (which we, without loss of generality, quantify with respect to the channel input). Bob then performs a homodyne detection to measure either x- or p-quadrature of the quantum state he receives. Alternatively, Bob can measure both quadratures simultaneously with a generally unbalanced heterodyne detector (double homodyne) with the balancing $t_{het}$ as shown in Fig. 1b. After a sufficient number of transmitted pulses are measured, trusted parties perform error correction and privacy amplification on their data to obtain the secret key. The lower bound on the secret key, secure against collective attacks (directly extendable to security against general attacks in the asymptotic regime [11]) is given by:
@@
Figure 1: a) Prepare and measure scheme of a CV QKD protocol with noisy squeezed states. Alice modulates the squeezed state, characterized by the squeezed variance $V$ and AS excess noise $\Delta V$, according to normally distributed random variables and sends the Gaussian modulated squeezed state through a quantum channel characterized by transmittance distribution $P(T)$ and excess noise $\epsilon$. All modulator imperfections are attributed to the source. Bob at his end performs homodyne detection or b) unbalanced heterodyne detection with $t_{het}$. c) Equivalent general source model, used for purification and security analysis, based on the balanced coupling of two oppositely squeezed states in modes $A_{1}$ and $B$ with different squeezed variances $V_{1}$, $V_{2}$ and a balanced heterodyne measurement at the Alice’s side, fed by a squeezed state with variance $V_{m}$ in mode $A_{2}$.
@@
The trusted parties may assume the AS noise $\Delta V$ to be either trusted (hence, being out of control by an eavesdropper) or untrusted (fully controlled by Eve). We further denote the trusted and untrusted AS noise as $\Delta V_{t}$ and $\Delta V_{u}$, respectively. We model the trusted AS noise by redefining the modulation variance of the AS quadrature $V_{p}\rightarrow V_{p}+\Delta V_{t}$ ($\Delta V_{t}$ does not contribute to the mutual information between Alice and Bob). On the other hand, the untrusted AS noise $\Delta V_{u}$ is added to the AS quadrature similarly to the channel noise $\epsilon$ (for a detailed explanation of the assumption that all noise is concentrated in the AS quadrature, see Sec.B).
@@
In this expression, the switching ratio between squeezed and AS quadrature measurements does not explicitly appear. This is because in the asymptotic regime, where the number of measurements approaches infinity, statistical fluctuations and biases due to finite sample sizes can be ignored. Consequently, the estimations of parameters like noise and channel transmission are exact, and no bias correction is necessary. Therefore, biases do not appear in the key rate expression for the asymptotic case.
However, in practical implementations with a finite number of signals, the switching ratio plays a crucial role. It affects the precision of parameter estimations and thus impacts the finite-size key rate. We will elaborate on the finite-size effects and the optimization of the switching ratio in subsequent sections (Sec.3.2).
@@
To study the optimal performance of CV QKD with untrusted AS noise (on top of symmetric noise $\epsilon$) in the asymptotic regime, we only consider the measurement of squeezed quadrature by Bob. A brief summary of all stages of the protocol implementation and relevant quantities are shown in Fig. 2.
Figure 2: Diagram of the protocol operation indicating relevant parameters and estimated quantities. State preparation optimization is based on estimated channel parameters and error correction efficiency $\beta$. Not trusting anti-squeezing noise will translate into significant increase of channel noise in respective quadrature. Monitoring is required to trust the anti-squeezing noise $\Delta V$. In a fiber link channel loss remains fixed and (aside from channel noise) only lower bound on transmittance $T^{low}$ needs to be estimated. In a free-space link statistical properties of probability distribution of transmittance, namely $\langle T\rangle$ and $\langle\sqrt{T}\rangle$, are required to correctly assess the secure key. Note that only optimization of data fraction can be done after the block measurement has already been commenced, other parameters are optimized before the protocol round begins, i.e. prior to the generation and transmission of quantum states.
@@ 3.1 Asymptotic key rate
In the presence of channel noise, no simple analytical expressions can be obtained for the Holevo bound, as we rely on the 3-mode purification described in Sec.2. The presence of trusted noise can improve tolerance to untrusted channel noise, as can be seen from the plots in Fig. 3, where we model the performance of the protocol assuming realistic error correction efficiency of $\beta=0.95$ [45] and optimizing the modulation variance. The improvement by the trusted AS noise is more evident in DR than in RR, being a case of the ”fighting noise with noise” method [23, 44]. Interestingly, a moderately squeezed source ($V=0.5$ SNU) in the presence of optimal trusted AS noise has a higher tolerance to channel noise than the protocol with a higher degree of pure signal squeezing of the source state for DR. On the other hand, when the AS noise cannot be trusted, the tolerance of the protocol to channel noise is significantly reduced, as can be seen from the plots in Fig. 4, as the presence of untrusted AS noise increases the Holevo bound. This emphasises the need to control and to estimate noise in the AS quadrature, which would otherwise lead to underestimation of the Holevo bound.
@@ 3.2 Finite-size key rate
A necessary requirement to verify the security of QKD is to estimate the channel parameters from which the Holevo bound can be found. Channel parameters can only be estimated within a certain confidence interval, which is inversely propositional to the number of states sent to Bob. Typically part of the received data is used for parameter estimation, and the rest is used for the key generation [46]. It has been also shown that all the measurements can be used for both parameter estimation and key generation by performing error correction before parameter estimation [47]. For completeness, we consider both cases, when part of the measurements are disclosed for parameter estimation (see C), and when all the measurements are used for parameter estimation and key generation.
@@
In order not to underestimate the information available to the eavesdropper, we use the lower bound for the confidence interval of channel transmittance $T^{low}$ and the upper bound for channel noise $V_{\epsilon}^{up}$. We consider $\mathbf{E}(T^{low})=T-6.5\Upsilon$ and $\mathbf{E}(V_{\epsilon}^{up})=V_{\epsilon}+6.5s$ corresponding to the estimation error probability of $10^{-10}$, $\sigma$ and $s$ are the standard deviations of the channel transmittance and noise, respectively. In order to effectively use the available resources, we extend the parameter estimation from [10, 46] to use the measurements of both quadratures to potentially make a better estimation of the channel parameters (refer to C for details). We obtain the following finite-size key rate:
@@
In Fig. 5, our analysis reveals distinct trends based on the degree of squeezing in the source state. Imbalanced heterodyne measurements demonstrate superior performance with small block sizes ($N=10^{6}$) for highly squeezed states (dashed plots), while biased homodyne detection becomes more favorable with larger block sizes ($N=10^{7}$), see Fig. 55(b), 5(d) and 5(f). Notably, biased homodyne measurements consistently outperform other strategies for moderately squeezed states. In scenarios marked by elevated trusted noise and reduced block lengths (Fig. 55(e)), imbalanced heterodyne measurement proves significantly more efficacious than biased homodyne measurement. Furthermore, we present key rates when the AS quadrature measurement in imbalanced heterodyne measurements is disclosed publicly (red plots in Fig. 5 and 6), showcasing its potential contributions to parameter estimation and the selection of optimal error correction codes.
Fig. 6 adds complexity to the narrative, revealing that the efficacy of imbalanced heterodyne measurements is contingent on the presence of channel noise. Specifically, imbalanced heterodyne measurement emerges as consistently superior in scenarios characterized by a substantial amount of channel noise ($\epsilon>0.1$ SNU), while biased homodyne measurement excels in low-channel-noise conditions ($\epsilon<0.1$ SNU). This is because, in the presence of substantial noise ($\epsilon>0.1SNU$), more measurements of both quadratures are required to reduce the variance of the noise estimator, as it is proportional to the noise itself. For heterodyne measurements, this is not an issue since the number of measurements for both quadratures is always equal. In contrast, for homodyne measurements, Bob must perform additional measurements of the AS quadrature to estimate the AS noise. These AS measurements do not contribute to the key and are solely used for estimating the AS noise. The interplay of measurement strategies with channel noise highlights the nuanced considerations necessary for optimizing squeezed-state CV quantum communication systems.